OpenStudy (anonymous):

Find the equation of the tangent line x^4 + xy + y^4 3 at (1,1)

OpenStudy (anonymous):

rewrite the equation clearly please.

OpenStudy (anonymous):

\[x ^{4}+xy+y ^{4}=3 \] at (1,1)

OpenStudy (anonymous):

the answer is f(x)=-1

OpenStudy (anonymous):

Can you elaborate more on the steps you took if you don't mind?

OpenStudy (anonymous):

Ok give me 5 min. i will explain. Do you know what does derivative mean? in geometric view?

OpenStudy (anonymous):

I know \[y-y _{1}=m(x-x _{1})\] is used right? I have a problem finding the slope.

OpenStudy (anonymous):

wait i think i made a mistake. -1 is just the slope.

OpenStudy (anonymous):

can you solve other part of the problem if i explain how i found the slope?

OpenStudy (anonymous):

there i attached it. take a look

OpenStudy (anonymous):

now that you know the slope you can use the formula you wrote below. :)

OpenStudy (anonymous):

anything wrong?

OpenStudy (anonymous):

y-1=(-1)(x-1)=>y-1=1-x => y=2-x thats what ur looking for i guess.

OpenStudy (anonymous):

Just comparing it to what I have.

OpenStudy (anonymous):

so after you get 5yprime + 5 = 0. how does yprime then become -1?

OpenStudy (anonymous):

5 y '+5=0 => 5y'=-5 => divide both side by 5 and you will get y'=-1

OpenStudy (anonymous):

Gotcha :)

OpenStudy (anonymous):

very good :)

OpenStudy (anonymous):

Ok man, I figured it out now. Thanks a lot!!!

OpenStudy (anonymous):

Your welcome. :)